OpenStudy (1dfanatic):

What is the equation for a line that passes through the points (4,-4) and (-8,26)?

OpenStudy (kayders1997):

Do you know where to start?

OpenStudy (1dfanatic):

i think make a graph??

OpenStudy (kayders1997):

You could or you could use the slope formula do you know what that is?

OpenStudy (1dfanatic):

um mx+b?

OpenStudy (kayders1997):

And does it matter what equation you have like what type?

OpenStudy (kayders1997):

That's the slope intercept formula the formula to find slope is the one I'm looking for

OpenStudy (1dfanatic):

Um I'm not sure, it gives me options if that answers your question lol

OpenStudy (1dfanatic):

and I don't remember it sorry

OpenStudy (1dfanatic):

y1-y2 or something like that?

OpenStudy (kayders1997):

Okay do you wanna put them in the chat? So than I'll know what they are looking for

OpenStudy (kayders1997):

Y2-y1 ------- X2-X1 the line is a division sign

OpenStudy (1dfanatic):

a) y=-5/2x-6 b) y=5/2x-6 c) y=-5/2x+6 d) y=5/2x+6

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

So how do you want to do you point like what one do you want to label one and which one do you want to label 2?

OpenStudy (kayders1997):

It doesn't matter which one you call point 1 and which one you call point 2 as long as your consistent with the formula

OpenStudy (1dfanatic):

I still dont get it im sorry

OpenStudy (kayders1997):

That's okay so let's say we call (4,-4) point 1 and (-8,26) point 2 is that okay?

OpenStudy (1dfanatic):

yes

OpenStudy (kayders1997):

Okay so point one you could call this (X1,y1) and since we called (4,-4) point one your X1 will be 4 and your y1 will be -4 understand? If not we can go to twiddle than you can see it on kinds of a whiteboard

OpenStudy (1dfanatic):

Okay i got that(:

OpenStudy (kayders1997):

Okay good! Now we do the same for the other part point 2 (x2,y2) since we called point 2 (-8,26) than x2 will be -8 and y2 will be 26

OpenStudy (kayders1997):

https://www.twiddla.com/2539472

OpenStudy (kayders1997):

Now it might get a little harder to show so if you can click on that link I could probably show it to you which will be much easier

OpenStudy (1dfanatic):

I clicked on it

OpenStudy (kayders1997):

Idk if you see a little chat box but I just wanna want sure your okay up to this point

OpenStudy (kayders1997):

?

OpenStudy (kayders1997):

We weren't done